A container of groceries has a mass of 15.0 kg. if its coefficent of static friction is 0.20 what force is needed to start the container to move?

PLEASE SHOW WORK
Answer #1

Force needed = uN = u*m*g = 0.2 * 15 * 9.81 = 29.43N (answer)